# A contramodule generalization of Neeman’s flat and projective module theorem

### v1

@@ Abstract.
This paper builds on top of [26]. We consider a complete, separated topological ring $\mathfrak{R}$ with a countable base of neighborhoods of zero consisting of open two-sided ideals. The main result is that the homotopy category of projective left $\mathfrak{R}$-contramodules is equivalent to the derived category of the exact category of flat left $\mathfrak{R}$-contramodules. In other words, a complex of flat $\mathfrak{R}$-contramodules is contraacyclic in the sense of Becker if and only if it is an acyclic complex with flat $\mathfrak{R}$-contramodules of cocycles.
@@ Introduction
The flat and projective periodicity theorem of Benson and Goodearl [6, Theorem 2.5] claims that, in the context of modules over an arbitrary ring $R$, if $0\longrightarrow F\longrightarrow P\longrightarrow F\longrightarrow 0$ is a short exact sequence with a flat module $F$ and a projective module $P$, then the module $F$ is actually projective. This result was rediscovered by Neeman [18, Remark 2.15], who proved the following much stronger claims.
@@
See [18, Theorem 8.6].
The aim of this paper is to prove the following generalization of Neeman’s theorem. Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Let $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of flat left $\mathfrak{R}$-contramodules. Then the following four conditions are equivalent:
@@
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ belongs to the closure of the class of all contractible complexes of flat $\mathfrak{R}$-contramodules under extensions and directed colimits;
@@
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is an $\aleph_{1}$-directed colimit of totalizations of short exact sequences of complexes of countably presentable flat $\mathfrak{R}$-contramodules.
@@
The equivalence of the latter three conditions (iic) $\Longleftrightarrow$ (iiic) $\Longleftrightarrow$ (ivc) was established in the paper [26, Theorem 13.2]. The aim of the present paper is to prove the equivalence of the former two conditions (ic) $\Longleftrightarrow$ (iic).
@@
The study of arbitrary (unbounded) complexes of projective modules/objects goes back to Jørgensen’s paper [12]. The terminology “contraderived category” was introduced by the present author in [19, 20], inspired by Keller’s terminology “coderived category” in [13] (see also [14]). The contemporary point of view on the topic came with Becker’s paper [4, Proposition 1.3.6(1)]. Let us spell out some definitions.
@@
Let $\mathsf{B}$ be an abelian category with enough projective objects. A complex $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$ is said to be contraacyclic (in the sense of Becker) if, for every complex of projective objects $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$, any morphism of complexes $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero. The quotient category $\mathsf{D}^{\mathsf{bctr}}(\mathsf{B})=\mathsf{Hot}(\mathsf{B})/\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})$ of the homotopy category $\mathsf{Hot}(\mathsf{B})$ of complexes in $\mathsf{B}$ by the thick subcategory of contraaacyclic complexes $\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})\subset\mathsf{Hot}(\mathsf{B})$ is called the (Becker) contraderived category of the abelian category $\mathsf{B}$. We refer to [24, Section 7] and [30, Remark 9.2] for a general discussion of the history and philosophy of the coderived and contraderived categories.
@@
For the abelian category of $R$-modules $\mathsf{B}=R{\operatorname{\mathsf{--Mod}}}$, the contraderived category is equivalent to the homotopy category of complexes of projective objects, $\mathsf{Hot}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{proj}})\simeq\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}})$[18, Proposition 8.1]. The equivalence of two conditions (i) and (ii) above (which is a part of [18, Theorem 8.6]) can be restated by saying that a complex of flat $R$-modules is contraacyclic if and only if it is acyclic as a complex in the exact category of flat $R$-modules $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$. It follows that the contraderived category of $R$-modules is equivalent to the derived category of the exact category of flat $R$-modules, $\mathsf{Hot}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{proj}})\simeq\mathsf{D}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}})\simeq\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}})$.
For the abelian category of $\mathfrak{R}$-contramodules $\mathsf{B}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$, the contraderived category is also equivalent to the homotopy category of complexes of projective objects, $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. This is a quite general result, valid for any locally presentable abelian category $\mathsf{B}$ with enough projective objects (which means, in particular, for the category of left contramodules over any complete, separated topological ring $\mathfrak{R}$ with a base of neighborhoods of zero formed by open right ideals) [30, Corollary 7.4]. When $\mathfrak{R}$ has a countable base of neighborhoods of zero consisting of open two-sided ideals, the results of the present paper allow us to say more.
@@
The equivalence of two conditions (ic) and (iic) above (which is claimed in our Theorem 7.1) can be restated by saying that a complex of flat $\mathfrak{R}$-contramodules is contraacyclic if and only if it is acyclic as a complex in the exact category of lfat $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$. It follows that the contraderived category of $\mathfrak{R}$-contramodules is equivalent to the derived category of the exact category of flat contramodules, $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\simeq\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$.
@@
We refer to the paper [32] for a discussion of another generalization of the results of Neeman’s paper [18], from modules to curved DG-modules. In the present paper, we restrict ourselves to a contramodule generalization of [18, Theorem 8.6].
@@
Let us explain the connection with the Benson–Goodearl periodity theorem. Let $0\longrightarrow F\longrightarrow P\longrightarrow F\longrightarrow 0$ be a short exact sequence of modules with a flat $R$-module $F$ and a projective $R$-module $P$. Splicing copies of this short exact sequence on top of each other in both the positive and negative cohomological directions, we obtain an unbounded complex of projective $R$-modules $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ with flat $R$-modules of cocycles. So the complex $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}=P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ satisfies (ii), and consequently it also satisfies (i). Thus the identity morphism of complexes $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow F^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ must be homotopic to zero. Hence the complex $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contractible, and it follows that its modules of cocycles are projective. This is the argument of [18, Remark 2.15].
@@
Conversely, given a complex of projective $R$-modules with flat $R$-modules of cocycles, one can chop it up into short exact sequence pieces, shift them to one and the same cohomological degree, and take the countable direct sum. This produces a short exact sequence of the form $0\longrightarrow F\longrightarrow P\longrightarrow F\longrightarrow 0$. This argument shows that the Benson–Goodearl periodicity theorem is essentially equivalent to the assertion that in any complex of projective modules with flat modules of cocycles, the modules of cocycles are actually projective [8, proof of Proposition 7.6], [11, Propositions 1 and 2], [2, Proposition 2.4].
@@
Similarly, one can start with the equivalence of two conditions (ic) and (iic) and deduce the following version of flat and projective periodicity theorem for contramodules. If $0\longrightarrow\mathfrak{F}\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{F}\longrightarrow 0$ is a short exact sequence of $\mathfrak{R}$-contramodules with a flat contramodule $\mathfrak{F}$ and a projective contramodule $\mathfrak{P}$, then the contramodule $\mathfrak{F}$ is actually projective. A direct argument deducing this assertion from the Benson–Goodearl periodicity theorem for modules was spelled out in [26, Proposition 12.1]. The equivalence of two conditions (ic) and (iic), proved in Theorem 7.1 in the present paper, is a much stronger result.
@@
We refer to the introduction to the paper [3] and to the preprint [27, Sections 7.8 and 7.10] for a further discussion of periodicity theorems. Before we finish this introduction, let us say a few words about contramodules.
@@
The categories of topological modules are usually not abelian; and indeed, contramodules are not topological modules. The $\mathfrak{R}$-contramodules are $\mathfrak{R}$-modules endowed with infinite summation operations with the coefficient families converging to zero in the topology of $\mathfrak{R}$[22, Section 2.1], [28, Sections 1.2 and 5], [29, Section 6]. There is a naturally induced topology on a left $\mathfrak{R}$-contramodule, but it need not be even separated. So $\mathfrak{R}$-contramodules are nontopological modules over a topological ring $\mathfrak{R}$, forming an abelian category (in fact, a locally presentable abelian category with enough projective objects; cf. [30, Sections 6–7]).
When the topological ring $\mathfrak{R}$ has a countable base of neighborhoods of zero, the contramodule theory simplifies quite a bit. In particular, all left $\mathfrak{R}$-contramodules are complete (though they still need not be separated) in their induced topology. Furthermore, there is a well-behaved full subcategory of flat $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}\subset\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Flat left $\mathfrak{R}$-contramodules are separated and complete in their induced topology. The full subcategory $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is resolving and closed under directed colimits in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Moreover, the directed colimits are exact functors in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$[28, Sections 6–7] (but not in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$[28, Examples 4.4]).

### v3

@@ Abstract.
This paper builds on top of [34]. We consider a complete, separated topological ring $\mathfrak{R}$ with a countable base of neighborhoods of zero consisting of open two-sided ideals. The main result is that the homotopy category of projective left $\mathfrak{R}$-contramodules is equivalent to the derived category of the exact category of flat left $\mathfrak{R}$-contramodules, and also to the homotopy category of flat cotorsion left $\mathfrak{R}$-contramodules. In other words, a complex of flat $\mathfrak{R}$-contramodules is contraacyclic (in the sense of Becker) if and only if it is an acyclic complex with flat $\mathfrak{R}$-contramodules of cocycles, and if and only if it is coacyclic as a complex in the exact category of flat $\mathfrak{R}$-contramodules. These are contramodule generalizations of theorems of Neeman and of Bazzoni, Cortés-Izurdiaga, and Estrada.
@@ Introduction
The flat and projective periodicity theorem of Benson and Goodearl [6, Theorem 2.5] claims that, in the context of modules over an arbitrary ring $R$, if $0\longrightarrow F\longrightarrow P\longrightarrow F\longrightarrow 0$ is a short exact sequence with a flat module $F$ and a projective module $P$, then the module $F$ is actually projective. This result was rediscovered by Neeman [26, Remark 2.15], who proved the following much stronger claims.
@@
See [26, Theorem 8.6].
A further development came in the paper by Bazzoni, Cortés-Izurdiaga, and Estrada [2], where the cotorsion periodicity theorem [2, Theorem 1.2(2), Proposition 4.8(2), or Theorem 5.1(2)] was proved. It claims that if $0\longrightarrow M\longrightarrow C\longrightarrow M\longrightarrow 0$ is a short exact sequence of $R$-modules and the $R$-module $C$ is cotorsion, then the $R$-module $M$ is cotorsion, too. As a corollary, it was essentially shown in [2, Theorem 5.3] that the following two conditions on a complex of flat left $R$-modules $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ are equivalent to (i–iii):
  1. (iv)
    for every complex of cotorsion left $R$-modules $C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $R$-modules $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero;
  2. (v)
    for every complex of flat cotorsion left $R$-modules $G^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $R$-modules $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow G^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero.
The aim of this paper is to prove the following generalization of the theorems of Neeman and of Bazzoni, Cortés-Izurdiaga, and Estrada. Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Let $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex of flat left $\mathfrak{R}$-contramodules. Then the following six conditions are equivalent:
@@
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ can be obtained from contractible complexes of flat $\mathfrak{R}$-contramodules using extensions and directed colimits;
@@
    $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is an $\aleph_{1}$-directed colimit of total complexes of short exact sequences of complexes of countably presentable flat $\mathfrak{R}$-contramodules;
  5. (vc)
    for every complex of cotorsion left $\mathfrak{R}$-contramodules $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero;
  6. (vic)
    for every complex of flat cotorsion left $\mathfrak{R}$-contramodules $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes of $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero.
@@
The equivalence of the three conditions (iic) $\Longleftrightarrow$ (iiic) $\Longleftrightarrow$ (ivc) was established in the paper [34, Theorem 13.2]. The aim of the present paper is to prove the equivalence of the four conditions (ic) $\Longleftrightarrow$ (iic) $\Longleftrightarrow$ (vc) $\Longleftrightarrow$ (vic).
@@
The study of arbitrary (unbounded) complexes of projective modules goes back to Jørgensen’s paper [19]. The terminology “contraderived category” was introduced by the present author in [27, 28], inspired by Keller’s terminology “coderived category” in [21] (see also [22]). The contemporary point of view on the topic came with Becker’s paper [4, Proposition 1.3.6]. Let us spell out some definitions.
@@
Let $\mathsf{B}$ be an abelian (or exact) category with enough projective objects. A complex $B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$ is said to be contraacyclic (in the sense of Becker) if, for every complex of projective objects $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{B}$, any morphism of complexes $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow B^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero. The triangulated Verdier quotient category $\mathsf{D}^{\mathsf{bctr}}(\mathsf{B})=\mathsf{Hot}(\mathsf{B})/\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})$ of the homotopy category $\mathsf{Hot}(\mathsf{B})$ of complexes in $\mathsf{B}$ by the thick subcategory of contraaacyclic complexes $\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})\subset\mathsf{Hot}(\mathsf{B})$ is called the (Becker) contraderived category of the abelian/exact category $\mathsf{B}$.
Dually, let $\mathsf{A}$ be an exact category with enough injective objects. A complex $A^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{A}$ is said to be coacyclic (in the sense of Becker) if, for every complex of injective objects $J^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{A}$, any morphism of complexes $A^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow J^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero. The quotient category $\mathsf{D}^{\mathsf{bco}}(\mathsf{A})=\mathsf{Hot}(\mathsf{A})/\mathsf{Ac}^{\mathsf{bco}}(\mathsf{A})$ of the homotopy category $\mathsf{Hot}(\mathsf{A})$ by the thick subcategory of coacyclic complexes $\mathsf{Ac}^{\mathsf{bco}}(\mathsf{A})\subset\mathsf{Hot}(\mathsf{A})$ is called the (Becker) coderived category of the exact category $\mathsf{A}$. We refer to [32, Section 7] and [40, Remark 9.2] for a general discussion of the history and philosophy of the coderived and contraderived categories. In the context of the present paper, the discussion in [36, Section 7] is relevant.
Generally speaking, for the abelian category $R{\operatorname{\mathsf{--Mod}}}$ of modules over a ring $R$, the coderived category $\mathsf{D}^{\mathsf{bco}}(R{\operatorname{\mathsf{--Mod}}})$ is quite different from the contraderived category $\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}})$. For the abelian category of $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$, the construction of the coderived category does not even make sense, as there are no nonzero injective objects in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ in general. However, the coderived category of the exact category of flat $R$-modules agrees with the derived and contraderived categories of the same exact category [36, Sections 7.4 and 7.6]. In this paper we generalize these results to the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ of flat contramodules over a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals.
@@
For the abelian category of $R$-modules $\mathsf{B}=R{\operatorname{\mathsf{--Mod}}}$, the contraderived category is equivalent to the homotopy category of complexes of projective objects, $\mathsf{Hot}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{proj}})\simeq\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}})$[26, Proposition 8.1]. The equivalence of two conditions (i) and (ii) above (which is a part of [26, Theorem 8.6]) can be restated by saying that a complex of flat $R$-modules is contraacyclic if and only if it is acyclic as a complex in the exact category of flat $R$-modules $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$. Besides, it is obvious that a complex in the exact category $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ is contraacyclic if and only if it is contraacyclic as a complex in $R{\operatorname{\mathsf{--Mod}}}$, because the projective objects of $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ coincide with those of $R{\operatorname{\mathsf{--Mod}}}$. It follows that the contraderived category of $R$-modules is equivalent to the derived category of the exact category of flat $R$-modules, $\mathsf{Hot}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{proj}})\simeq\mathsf{D}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}})=\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}})\simeq\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}})$.
@@
On the other hand, the injective objects of the exact category $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ are the flat cotorsion $R$-modules. So the equivalence of two conditions (ii) and (v) above (which follows immediately from [2, Theorem 5.3]) can be restated by saying that the classes of acyclic and coacyclic complexes in the exact category $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ coincide. It also follows from [2, Theorem 5.3] that the coderived category of the exact category $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ is equivalent to the homotopy category of complexes of flat cotorsion $R$-modules, $\mathsf{Hot}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bco}}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}})$. Thus we have $\mathsf{Hot}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{proj}})\simeq\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}})=\mathsf{D}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}})=\mathsf{D}^{\mathsf{bco}}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}})\simeq\mathsf{Hot}(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}^{\mathsf{cot}})$.
For the abelian category of $\mathfrak{R}$-contramodules $\mathsf{B}=\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$, the contraderived category is also equivalent to the homotopy category of complexes of projective objects, $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. This is a quite general result, valid for any locally presentable abelian category $\mathsf{B}$ with enough projective objects (which means, in particular, for the category of left contramodules over any complete, separated topological ring $\mathfrak{R}$ with a base of neighborhoods of zero formed by open right ideals) [40, Corollary 7.4]. When $\mathfrak{R}$ has a countable base of neighborhoods of zero consisting of open two-sided ideals, the results of the present paper allow us to say more.
The equivalence of two conditions (ic) and (iic) above (which is claimed in our Theorem 11.1) can be restated by saying that a complex of flat $\mathfrak{R}$-contramodules is contraacyclic if and only if it is acyclic as a complex in the exact category of flat $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$. Similarly to the module case, it is obvious that a complex in the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is contraacyclic if and only if it is contraacyclic as a complex in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$, because the projective objects in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ coincide. It follows that the contraderived category of $\mathfrak{R}$-contramodules is equivalent to the derived category of the exact category of flat contramodules, $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\simeq\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$.
The equivalence of two conditions (iic) and (vic) above (which is another assertion of our Theorem 11.1) can be restated by saying that a complex of flat $\mathfrak{R}$-contramodules is coacyclic in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ if and only if it is acyclic in this exact category. Furthermore, according to our Corollary 9.2, the coderived category of the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is equivalent to the homotopy category of complexes of flat cotorsion $\mathfrak{R}$-contramodules, $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bco}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$. So we have $\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{proj}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})=\mathsf{D}^{\mathsf{bco}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\simeq\mathsf{Hot}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}^{\mathsf{cot}})$, just as in the module case.
@@
We refer to the paper [43] for a discussion of another generalization of the results of Neeman’s paper [26] and the paper of Bazzoni, Cortés-Izurdiaga, and Estrada [2], from modules to curved DG-modules. In the present paper, we discuss contramodule generalizations of [26, Theorem 8.6] and [2, Theorem 5.3].
@@
Let us explain the connection with the Benson–Goodearl periodity theorem. Let $0\longrightarrow F\longrightarrow P\longrightarrow F\longrightarrow 0$ be a short exact sequence of modules with a flat $R$-module $F$ and a projective $R$-module $P$. Splicing copies of this short exact sequence on top of each other in both the positive and negative cohomological directions, we obtain an unbounded acyclic complex of projective $R$-modules $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ with flat $R$-modules of cocycles. So the complex $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}=P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ satisfies (ii), and consequently it also satisfies (i). Thus the identity morphism of complexes $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow F^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ must be homotopic to zero. Hence the complex $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contractible, and it follows that its modules of cocycles are projective. This is the argument of [26, Remark 2.15].
@@
Conversely, given an acyclic complex of projective $R$-modules with flat $R$-modules of cocycles, one can chop it up into short exact sequence pieces, shift them to one and the same cohomological degree, and take the countable direct sum. This produces a short exact sequence of the form $0\longrightarrow F\longrightarrow P\longrightarrow F\longrightarrow 0$. This argument shows that the Benson–Goodearl periodicity theorem is essentially equivalent to the assertion that in any acyclic complex of projective modules with flat modules of cocycles, the modules of cocycles are actually projective [10, proof of Proposition 7.6], [13, Propositions 1 and 2], [2, Proposition 2.4].
@@
Similarly, one can start with the equivalence of two conditions (ic) and (iic) and deduce the following version of flat and projective periodicity theorem for contramodules. If $0\longrightarrow\mathfrak{F}\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{F}\longrightarrow 0$ is a short exact sequence of $\mathfrak{R}$-contramodules with a flat contramodule $\mathfrak{F}$ and a projective contramodule $\mathfrak{P}$, then the contramodule $\mathfrak{F}$ is actually projective. A direct argument deducing this assertion from the Benson–Goodearl periodicity theorem for modules was spelled out in [34, Proposition 12.1]. The equivalence of two conditions (ic) and (iic), proved in Theorem 11.1 in the present paper, is a much stronger result.
Analogous arguments are applicable to the cotorsion periodicity of Bazzoni, Cortés-Izurdiaga, and Estrada. The cotorsion periodicity theorem for $R$-modules, stated above in this introduction, is essentially equivalent to the assertion that in any acyclic complex of cotorsion $R$-modules, the $R$-modules of cocycles are actually cotorsion  [2, Theorem 1.2(2), Proposition 4.8(2), and Theorem 5.1(2)]. Similarly, the cotorsion periodicity theorem for $\mathfrak{R}$-contramodules claims that if $0\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{C}\longrightarrow\mathfrak{M}\longrightarrow 0$ is a short exact sequence of $\mathfrak{R}$-contramodules with a cotorsion contramodule $\mathfrak{C}$, then $\mathfrak{M}$ is also a cotorsion contramodule [34, Theorem 12.3]. This theorem can be restated by saying that in any acyclic complex of cotorsion $\mathfrak{R}$-contramodules $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, the contramodules of cocycles are cotorsion [34, Corollary 12.4].
@@
We refer to the introduction to the paper [3] and to the preprint [36, Sections 7.8 and 7.10] for a further discussion of periodicity theorems. Before we finish this introduction, let us say a few words about contramodules.
@@
The categories of topological modules are usually not abelian; and indeed, contramodules are not topological modules. The $\mathfrak{R}$-contramodules are $\mathfrak{R}$-modules endowed with infinite summation operations with the coefficient families converging to zero in the topology of $\mathfrak{R}$[30, Section 2.1], [38, Sections 1.2 and 5], [39, Section 6]. There is a naturally induced topology on a left $\mathfrak{R}$-contramodule, but it need not be even separated. So $\mathfrak{R}$-contramodules are nontopological modules over a topological ring $\mathfrak{R}$, forming an abelian category (in fact, a locally presentable abelian category with enough projective objects; cf. [40, Sections 6–7]).
When the topological ring $\mathfrak{R}$ has a countable base of neighborhoods of zero, the contramodule theory simplifies quite a bit. In particular, all left $\mathfrak{R}$-contramodules are complete (though they still need not be separated) in their induced topologies. Furthermore, there is a well-behaved full subcategory of flat $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}\subset\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Flat left $\mathfrak{R}$-contramodules are separated and complete in their induced topologies. The full subcategory $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is resolving and closed under directed colimits in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. Moreover, the directed colimits are exact functors in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$[38, Sections 6–7] (but not in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$[38, Examples 4.4]).